# VTAO-BiManip: Masked Visual-Tactile-Action Pre-training with Object Understanding for Bimanual Dexterous Manipulation

### v1

@@ III-B1 VTAO Encoder $E_{\theta}(V,C,A)\rightarrow(h_{\text{cls}},h_{v},h_{c},h_{a},\allowbreak h_{o})$
To enhance the feature extraction ability of the VTAO encoder $E_{\theta}$ for downstream tasks, we define four reconstruction objectives for the decoder: a) reconstruct the image $\widehat{V}$; b) reconstruct tactile signals $\widehat{C}$; c) reconstruct current actions and predict future actions $[\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}]$; d) reconstruct the rotation and size information of the bottle $\widehat{O}$, as shown in Fig. 3 (left). Here, $m$ denotes the mask tokens used for feature alignment.
Reconstruct Image $\widehat{V}$ and Tactile Signals $\widehat{C}$
@@
Building on our previous work [6], we employ a Transformer Decoder and MLP layers to reconstruct the image $\widehat{V}$ and tactile signals $\widehat{C}$ from the corresponding tokens $h_{v}$, $h_{c}$, respectively. Serving as an auxiliary task, this approach has been demonstrated to benefit downstream applications.
Reconstruct Current Actions and Predict Future Actions $[\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}]$
@@
To enable the encoder $E_{\theta}$ to acquire relevant prior knowledge of human manipulation tasks from single-frame VTAO input, we use a Transformer decoder and an MLP layer to reconstruct the dual-hand action of current step $\widehat{A}$ and the action sequences for the next p steps $[\widehat{A_{+1}},...,\widehat{A_{+p}}]$ from the corresponding action tokens $h_{a}$.
@@
Knowing the pose and size of the object is beneficial for downstream tasks such as bottle cap twisting. We enable the network to reconstruct the 6-DoF pose of the bottle relative to the camera coordinate system, along with the dimensions of the bottle cap and body, from object tokens $h_{\text{o}}$. This enables the encoder $E_{\theta}$ to better focus on the relevant information of the manipulated object.
@@
$L(\theta)=$ $W_{\text{img}}\cdot L(V,\widehat{V})+W_{\text{tac}}\cdot L(C,\widehat{C})+W_{\text{bot}}\cdot L(O,\widehat{O})$
$W_{\text{act}}\cdot L([A,A_{+1},...,A_{+p}],[\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}]),$ (2)
where $W_{\text{img}},W_{\text{tac}},W_{\text{bot}}$, and $W_{\text{act}}$ are weights and set to 1, 2, 5, 2, respectively. $L(\cdot)$ denotes the Euclidean distance between two vectors.
III-C VTAO RL for bimanual manipulations
III-C1 Problem Formulation
@@
III-C2 Task
To assess the effectiveness of VTAO pre-training on downstream tasks, we implement a bimanual bottle-cap unscrewing task in Isaac gym [49]. The setup positions the bottle body at the center of the table, with the left hand to its left in a pre-grasp gesture and the right hand fixed above it, as shown in the submitted video. The bottle body and cap are linked by a joint that rotates around the z-axis.
III-C3 State Space $\mathcal{S}$
We define the state space as $\mathcal{S}=\{V_{\text{obs}},C_{\text{obs}},\mathcal{P}\}$, where $V_{\text{obs}}$ is the ego-centric RGB image, $C_{\text{obs}}$ is the binary tactile signal from both hands, thresholded by $\lambda$, and $\mathcal{P}$ is proprioception data of both hands, including joint angles and velocities. To maintain feature representation from pre-training and improve stability in RL with sparse rewards, we freeze the parameters of the pre-trained VTAO model during RL training. The final feature vector input to the policy network is $\{h_{\text{cls}},\phi(\mathcal{P})\}$, where $h_{\text{cls}}=\mathrm{E}_{\theta_{f}}(V_{\text{obs}},C_{\text{obs}},A_{\text{obs}})$, with $\mathrm{E}_{\theta_{f}}$ representing the pre-trained VTAO encoder with frozen parameters. $A_{\text{obs}}\in\mathcal{P}$ represents the joint angles of both hands. $h_{\text{cls}}$ is the CLS token output by $\mathrm{E}_{\theta_{f}}$. $\phi(\cdot)$ is a linear layer.
@@ III-C4 Action Space $\mathcal{A}$
In the simulation, we use Shadow Hand [1] as the manipulator. This five-fingered robotic hand features 24 DoFs. Excluding the thumb, the distal joints of the four fingers are tendon-driven, resulting in a total of 20 DoFs. We implement dual-hand manipulation, where the base link of the right hand is fixed and the left hand can move freely with six DoFs. Consequently, the action $a=\pi_{\theta}(s)\in\mathbb{R}^{46}$.
III-C5 Learning Curriculum
It is challenging to train precise dual-hand manipulation with non-fixed objects from scratch. To address this, we use a two-stage learning curriculum to stabilize the learning process. In the first stage, we fix the bottle body on the table, enabling the left hand to learn to grasp the bottle and the right hand to learn to unscrew the cap. In the second stage, we release the bottle body, allowing both hands to collaborate in unscrewing the cap.
III-C6 Reward Function $\mathcal{R}$
To coordinate the dual dexterous hands in complex action spaces across different training stages, our reward function $r_{n}(n\in\{1,2\})$ is divided into two phases. In each phase, the total reward is the sum of the individual rewards for the left and right hands, $r_{n}=r_{\text{left}_{n}}+r_{\text{right}_{n}}$. The reward function of the first stage $r_{1}$ is:
{$r_{\text{left}_{1}}&=r_{\text{hdis}}+r_{\text{fcon}}\\ r_{\text{right}_{1}}&=r_{\text{cang}}+r_{\text{cvel}}+r_{\text{fdis}}\end{aligned}\right..$ (3)
@@ III-C6 Reward Function $\mathcal{R}$
Here, $r_{\text{hdis}}=\alpha_{1}\cdot d_{\text{h2b}}$ is a distance penalty, where $d_{\text{h2b}}$ is the distance between the left hand and the bottle. $r_{\text{fcon}}=\alpha_{2}\cdot N_{\text{con}}$ incentivizes multiple contacts between the left hand and the bottle, with $N_{\text{con}}$ representing the number of contacts. The rotation angle reward, $r_{\text{cang}}=\alpha_{3}\cdot\min(a_{\text{c}},7.0)$, depends on the rotation angle of the bottle cap $a_{\text{c}}$. $r_{\text{cvel}}=\alpha_{4}\cdot C_{\text{flag}}\cdot v_{\text{c}}$ rewards rotation speed, where $v_{\text{c}}$ is the rotation speed of the cap, and $C_{\text{flag}}$ equals 1 if the right hand is in contact with the cap, and 0 otherwise. The term $r_{\text{fdis}}=\alpha_{5}\cdot\exp(-10\cdot d_{\text{fz}})$ encourages the right-hand fingertips to approach the bottle cap, with $d_{\text{fz}}$ being the z-axis distance between the fingertips and the height of the cap. The weights $\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$, $\alpha_{4}$, and $\alpha_{5}$ are set to $-5$, $0.05$, $0.5$, $1.1$, and $0.5$, respectively.
The rewards for the second phase, $r_{2}$ are as follows:
{$r_{\text{left}_{2}}&=r_{\text{htdis}}+r_{\text{bdis}}+r_{\text{brot}}\\ r_{\text{right}_{2}}&=r_{\text{right}_{1}}\end{aligned}\right.,$ (4)
@@
where $r_{\text{htdis}}=\beta_{1}\cdot\exp(-5\cdot d_{\text{h2t}})$ is the approaching reward, with $d_{\text{h2t}}$ representing the distance between the left hand and the target position. $r_{\text{bdis}}=\beta_{2}\cdot\exp(-10\cdot d_{\text{o2i}})$ encourages the bottle to remain near its initial position, with $d_{\text{o2i}}$ denoting the distance between the bottle’s current and initial position. The reward $r_{\text{brot}}=\beta_{3}\left(\frac{1}{|d_{\text{qua}}|+1}\right)$ encourages the axis of the bottle remains perpendicular to the table, where $d_{\text{qua}}=2\cdot\arcsin\left(\min\left(\lVert q_{\text{diff}}\rVert_{2},1.0\right)\right)$ , and $q_{\text{diff}}=q_{\text{bot}}\cdot q_{\text{ini}}$ represents the vector part of the quaternion difference between the bottle’s current quaternion $q_{\text{bot}}$ and its initial orientation $q_{\text{ini}}$. The weights $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$ are all set to 1.0.
@@ IV-A Experiment Settings
IV-A1 Dataset
Using our VTAO acquisition system, we collect 216 human bimanual manipulation trajectories involving 26 different bottles. Each trajectory lasts 6 to 17 seconds and comprises a total of 61684 frames. Each frame contains aligned visual, tactile, and action data, along with bottle pose and size labels, as shown in the submitted video. Before training VTAO-BiManip, we preprocess the data by cropping images to $224\times 224$ pixels, binarizing tactile data with a $\lambda=0.4V$ threshold, aligning bottle pose with the camera coordinate system of the visual data, and retargeting human hand movements to the joint angles of the dual dexterous hands. The final data representations include images $V\in\mathbb{R}^{224\times 224\times 3}$, tactile data $C\in\mathbb{R}^{40}$, and action data $A\in\mathbb{R}^{48}$.
IV-A2 Metrics
We use success rate to evaluate the effectiveness of our method on downstream tasks. A task is considered successful if the dexterous hand rotates the bottle cap by more than half a turn. During training, we calculate the success rate as the average of the last ten manipulations in each parallel environment. For evaluation, we select the model from the 3500th iteration of the second training phase and measure the average success rate across 10 manipulations for each bottle in the evaluation set.
IV-A3 Implementation Details
We select 15 bottles from ShapeNet [50] with varying body and cap sizes, using 10 for training (Seen) and 5 for testing (Unseen). Our VTAO-BiManip runs on a system with an Intel Xeon Gold 6326 and an NVIDIA 3090. During pretraining, we use the AdamW optimizer with a learning rate of 2e-5. The mask ratios are set to $r_{v}=0.75$, $r_{c}=0.5$, and $r_{a}=0.5$. Training with a minibatch size of 8 takes about 2 hours for 300 epochs. For RL, we run 400 parallel environments and train using the PPO algorithm for roughly 4500 iterations, requiring about 62 hours to train the dual-hand manipulation strategy network.
IV-A4 Baselines
To validate the effectiveness of our method, we design several baselines with different pre-training structures, as shown in Table I. We designate our baselines in the form of VTAO. Here, V, T, and A correspond to the pretraining of visual, tactile, and action modalities, respectively, while O represents the reconstruction of bottle information. Both tactile and action modalities involve data from two hands. The action modality entails reconstructing the current frame and predicting the next five frames. The absence of a letter in the naming indicates that the corresponding modality is not used. For methods that do not involve pretraining, we denote them with the suffix ‘-Scr’.
@@ IV-B Effectiveness of pre-training with different modalities
Figure 4: Qualitative results of different pre-train methods. The left is the training process; the right is the evaluation results
@@
By comparing VT, VTA, VTO, and VTAO, we find that incorporating action modality prediction and object understanding into visuo-tactile fusion training enhances downstream task success rates. This validates the effectiveness and complementarity of the action prediction task and the object understanding task proposed in our approach.
Both Base and A, which only possess proprioception of the robotic hands without environmental perception, fail in training. In contrast, the pretraining method using only tactile perception, T, demonstrates a certain level of manipulation capability, effectively demonstrating the ability of tactile perception to gather environmental information and its effectiveness in enhancing operational performance.
@@ IV-C1 Effectiveness of Bimanual Pre-training
Ours dual dual 5steps $24\times 2$checkmark 63±4 51±7

### v2

@@ III-B1 VTAO Encoder $E_{\theta}(V,T,A)\rightarrow(h_{\text{fs}},h_{v},h_{c},h_{a},\allowbreak h_{o})$
To enhance the feature extraction ability of the VTAO encoder $E_{\theta}$ for downstream tasks, we define four reconstruction objectives for the decoder: a) reconstruct the image $\widehat{V}$; b) reconstruct tactile signals $\widehat{T}$; c) reconstruct current actions $\widehat{A}$ and predict future actions $\widehat{A_{+1}},...,\widehat{A_{+p}}$] for $p$ steps; d) reconstruct the 6-DoF pose and 3-D size information of the bottle $\widehat{O}$, as shown in Fig. 3 (left). Here, $m$ denotes the mask tokens used to fill in missing dimensions of the input sequence.
Reconstruct Image $\widehat{V}$ and Tactile Signals $\widehat{T}$
@@
We employ a Transformer decoder and MLP layers to reconstruct the image $\widehat{V}$ and tactile signals $\widehat{T}$ from the corresponding tokens $h_{v}$ and $h_{c}$, respectively. This approach has been demonstrated to benefit downstream applications in [liu2024m2vtp].
Reconstruct Current Actions and Predict Future Actions $\widehat{A},\widehat{A_{+1}},...,\widehat{A_{+p}}$]
@@
To enable the encoder $E_{\theta}$ to acquire prior knowledge of human manipulation tasks from single-frame VTAO input, we use a Transformer decoder and an MLP layer to reconstruct the current dual-hand action $\widehat{A}$ and predict future actions $\widehat{A_{+1}},...,\widehat{A_{+p}}$] for p steps, based on the action tokens $h_{a}$.
@@
Our framework reconstructs the 6-DoF pose of the assembled object (bottle with an attached cap) relative to the camera’s 3D coordinate frame, along with its 3D bounding box dimensions (width, height, depth). This is achieved by decoding the object tokens $h_{\text{o}}$, which forces the encoder $E_{\theta}$ to learn geometry-dynamic representations. These representations encode structural priors of the object and its spatiotemporal state transitions, thereby enabling real-time motion adaptation during dynamic manipulation.
@@
[…]
where $W_{\text{img}},W_{\text{tac}},W_{\text{bot}}$, and $W_{\text{act}}$ are weights for each loss term, set to 1, 2, 5, and 2, respectively. $L(\cdot)$ denotes the Euclidean distance between two vectors.
III-C Curriculum RL for Dexterous Bimanual Manipulation
III-C1 Task
To assess the effectiveness of VTAO pre-training on downstream tasks, we implement a bimanual bottle-cap rotation task (focusing solely on the unscrewing process). The setup positions the bottle body at the center of the table with the left hand in a pre-grasp position (left side) and the right hand positioned above it, as shown in the submitted video. The bottle body and cap remain connected through a z-axis rotational joint. Successful operation is defined by cap rotation exceeding 4 rad (angular displacement) rather than physical removal, as defined in [liu2024m2vtp].
III-C2 Problem Formulation
@@
III-C3 State Space with Multi-Modal Pre-trained Representations $\mathcal{S}$
We define the state space as $\mathcal{S}=\{V_{\text{obs}},T_{\text{obs}},\mathcal{P}\}$, where $V_{\text{obs}}$ is the ego-centric RGB image, $T_{\text{obs}}$ is the binary tactile signal from both hands, thresholded by $\lambda$, and $\mathcal{P}$ is proprioception data of both hands, including joint angles and velocities. Notably, object state information is excluded from the input state space $\mathcal{S}$ during RL training, as such privileged data is difficult to directly measure under real-world deployment constraints. To maintain feature representation from pre-training and improve stability in RL with sparse rewards, we freeze the parameters of the pre-trained VTAO model during RL training. The final feature vector input to the policy network is $\{h_{\text{fs}},\phi(\mathcal{P})\}$, where $h_{\text{fs}}=\mathrm{E}_{\theta_{f}}(V_{\text{obs}},T_{\text{obs}},A_{\text{obs}})$, with $\mathrm{E}_{\theta_{f}}$ representing the pre-trained VTAO encoder with frozen parameters. $A_{\text{obs}}\in\mathcal{P}$ represents the joint angles of both hands. $h_{\text{fs}}$ is the FS token output by $\mathrm{E}_{\theta_{f}}$. $\phi(\cdot)$ is a linear layer.
@@ III-C4 Action Space $\mathcal{A}$
In the simulation, we use Shadow Hand [shadowhand] as the manipulator. This five-fingered robotic hand features 24 DoFs. Excluding the thumb, the distal joints of the four fingers are tendon-driven, resulting in a total of 20 DoFs. We implement dual-hand manipulation, where the right hand is controlled with its initial position serving as the reference point, while the left hand is freely manipulated with six DoFs. Consequently, the action $a=\pi_{\theta}(s)\in\mathbb{R}^{46}$.
III-C5 Learning Curriculum with Stage-specific Reward Design $\mathcal{R}$
It is challenging to train precise dual-hand manipulation with non-fixed objects from scratch. In our experiments, single-stage training yields a consistent zero success rate due to insufficient exploration in the high-dimensional bimanual action space. To overcome this and stabilize the learning process, we implement a progressive two-stage curriculum with phase-dependent reward functions $r_{n}(n\in\{1,2\})$, where each stage’s total reward combines the individual rewards of the left and right hands: $r_{n}=r_{\text{left}_{n}}+r_{\text{right}_{n}}$.
Stage 1: Sub-skill Learning. In the first stage, we fix the bottle body on the table, enabling the left hand to learn to grasp the bottle and the right hand to learn to unscrew the cap. The reward function of the first stage $r_{1}$ is:
  • Left hand learns bottle grasping through:
    $r_{\text{left}_{1}}=\underbrace{\alpha_{1}d_{\text{h2b}}}_{\text{distance penalty}}+\underbrace{\alpha_{2}N_{\text{con}}}_{\text{contact reward}}.$ (3)
    Here, $\alpha_{1}\!=\!-5$, $\alpha_{2}\!=\!0.05$. $d_{\text{h2b}}$ is the distance between the left hand and the bottle. $N_{\text{con}}$ represents the number of contacts between the left hand and the bottle body.
  • Right hand develops cap rotation skills via:
    $r_{\text{right}_{1}}=\underbrace{\alpha_{3}\min(a_{c},4.0)}_{\text{rotation angle reward}}+\underbrace{\alpha_{4}C_{\text{flag}}v_{c}}_{\text{rotation speed reward}}+\underbrace{\alpha_{5}e^{-10d_{\text{fz}}}}_{\text{approach reward}}.$ (4)
@@ III-C5 Learning Curriculum with Stage-specific Reward Design $\mathcal{R}$
    Here, $\alpha_{3}\!=\!0.5$, $\alpha_{4}\!=\!1.1$, $\alpha_{5}\!=\!0.5$. $a_{\text{c}}$ is the rotation angle of the bottle cap. $C_{\text{flag}}$ equals 1 if the right hand is in contact with the cap, and 0 otherwise. $v_{\text{c}}$ is the rotation speed of the cap. The approach reward term encourages the right-hand fingertips to approach the bottle cap, with $d_{\text{fz}}$ denoting the z-axis distance between the fingertips and cap’s top surface.
Stage 2: Collaborative Learning. In the second stage, we release the bottle body, allowing both hands to collaborate in unscrewing the cap. The rewards for the second stage, $r_{2}$ are as follows:
  • Left hand manages grasping and stabilization:
    $r_{\text{left}_{2}}=\underbrace{\beta_{1}e^{-5d_{\text{h2t}}}}_{\text{hand approach}}+\underbrace{\beta_{2}e^{-10d_{\text{o2i}}}}_{\text{bottle stability}}+\underbrace{\frac{\beta_{3}}{|d_{\text{qua}}|+1}}_{\text{upright maintenance}}.$ (5)
@@
    Here, $\beta_{1}\!=\!\beta_{2}\!=\!\beta_{3}\!=\!1.0$. $d_{\text{h2t}}$ is the distance between the left hand and the target position $p_{t}=(x_{0},y_{0}-0.06,z_{0})$, with $(x_{0},y_{0},z_{0})$ being the hand’s initial position. The bottle stability term encourages the bottle to remain near its initial position, with $d_{\text{o2i}}$ denoting the distance between the bottle’s current and initial position. The upright maintenance term encourages the axis of the bottle remains perpendicular to the table, where $d_{\text{qua}}=2\cdot\arcsin\left(\min\left(\lVert q_{\text{diff}}\rVert_{2},1.0\right)\right)$, and $q_{\text{diff}}=q_{\text{bot}}\cdot q_{\text{ini}}$ represents the vector part of the quaternion difference between the bottle’s current quaternion $q_{\text{bot}}$ and its initial orientation $q_{\text{ini}}$.
  • Right hand maintains stage 1 rewards $r_{\text{right}_{2}}=r_{\text{right}_{1}}$ while adapting to dynamic bottle movements.
@@ IV-A Experiment Settings
IV-A1 Metrics
We use success rate to evaluate the effectiveness of our method on downstream tasks. A task is considered successful only if the dexterous hand rotates the bottle cap by more than 4 rad. During training, we compute the success rate for each iteration as the moving average over the preceding ten episodes across all parallel environments. For evaluation, we select the model from the 3500th iteration of the second training stage and measure the average success rate across 10 manipulations for each bottle in the evaluation set.
IV-A2 Implementation Details
We deploy our bimanual bottle cap twisting task in the Isaac Gym simulation environment [makoviychuk2021isaac]. We select 15 bottles from ShapeNet [chang2015shapenet] with varying body and cap sizes, using 10 for training (Seen) and 5 for testing (Unseen). Our VTAO-BiManip runs on a system with an Intel Xeon Gold 6326 and an NVIDIA 3090. During pretraining, we use the AdamW optimizer with a learning rate of 2e-5. The mask ratios are set to $r_{v}=0.75$, $r_{c}=0.5$, and $r_{a}=0.5$. Training with a minibatch size of 8 takes about 2 hours for 300 epochs. For RL, the two-stage curriculum learning takes approximately 62 hours to learn the dual-hand manipulation strategy. In the first stage, we train the policy using PPO for 1000 iterations. The resulting model is then used as the initialization for the second stage, where we further train for 3500 iterations.
IV-A3 Baselines
To validate the effectiveness of our method, we design multiple baseline models with different configurations, as summarized in Table 1. The baselines follow a naming convention of VTAO, where each capital letter indicates the inclusion of a specific pretrained modality using the MAE architecture shown in Figure 3: V-Visual (RGB observations), T-Bimanual binary tactile signals, A-Bimanual action joints (current reconstruction + 5-frame prediction) and O - Bottle information reconstruction. The absence of a letter denotes exclusion of the corresponding modality. We also construct a baseline that trains the multi-modal representation and the action policy from scratch by end-to-end reinforcement learning VTA-Scr. The baseline employs a ResNet encoder for visual inputs and MLP networks for processing tactile and action modalities.
@@ IV-B Effectiveness of pre-training with different modalities
Figure 4: Qualitative results of different pre-train methods in simulation. Left: training process; Right: evaluation results.
@@
Analysis of modality integration reveals two key findings: First, the performance improvements from VT to VTA (incorporating action modality prediction) and to VTO (incorporating object understanding) demonstrate that both action prediction and object understanding significantly enhance visuo-tactile fusion. Furthermore, VTAO achieves the highest success rates by combining both modalities, indicating complementary benefits between action prediction and object understanding tasks.
Experiments with proprioception-only configurations (Base and A) fail to train, whereas Base+Obj achieves successful training. This underscores the critical importance of object state perception for enabling policy networks to dynamically adjust manipulation strategies based on target object status.
Although our method achieves the highest success rate, the overall performance (around 60%) remains relatively low, which we attribute to the task’s high complexity — specifically, the large variations in object shapes and sizes in our experimental setting and the high-dimensional action space of dual-hand manipulation.
@@ IV-C Ablation Study
We conduct comprehensive ablation studies to validate key design choices in our approach. All experimental results are summarized in the Fig. 5.
@@ IV-C1 Effectiveness of Bimanual Pre-training
Ours dual dual 5steps $24\times 2$ $\checkmark$ 63±4 51±7